Consider the market for kiwis in the small open economy of Isoland. This market is characterized by the following,
where Q is quantity of kiwis, and P is the price of a unit of kiwis in dollars:
Domestic Demand: P = 1,000 – 2Q
Domestic Supply: P = (1/2)Q
World Price = $100
Suppose now that Iceland's dictator decides to charge a $5 tariff per pound on imported kiwis. What is the
government’s tariff revenue with this tariff?
a. $2,800
b. $4,500
c. $4,800
d. $6,250
1
Expert's answer
2020-12-21T10:57:58-0500

Solution:

DD or QD:

P = 1000 - 2Q

100 = 1000 - 2Q

2Q = 1000 - 100

2Q = 900

Q = 450

Quantity Demanded = 450 kiwis


DS or QS:

P = 1/2Q

100 = 1/2Q

Q = 200

Quantity Supplied = 200 kiwis


Quantity of Kiwis imported = QD - QS

=450 - 200 = 250 kiwis



Revenue tariff = Tariff per unit * Units imported

= $5 * 250

=$1.250